What is ΔrH for a reaction that has ΔrG = -0.2 kJ mol-1 and ΔrS = 65.4 J mol-1 K-1 at 48.2 °C?

Express your answer in kJ mol-1 .
Answer #1

ΔGo = -0.2 KJ/mol

ΔSo = 65.4 J/mol.K

= 0.0654 KJ/mol.K

T= 48.2 oC

= (48.2+273) K

= 321.2 K

use:

ΔGo = ΔHo - T*ΔSo

-0.2 = ΔHo - 321.2 * 0.0654

ΔHo = 20.8 KJ/mol

Answer: 20.8 KJ/mol